# Analytic action principle in ${\cal N}=2$ AdS4 harmonic superspace

### v1

@@ 5 Example: component structure of the free hypermultiplet action
$D^{++}F^{+}=$ $0\,$, (5.8a)
$D^{++}M^{-}-({\cal J}+2S^{--}D^{++})F^{+}=$ $0\,$, (5.8b)
$D^{++}N^{-}-({\cal J}+2S^{--}D^{++})F^{+}=$ $0\,$, (5.8c)
$D^{++}A^{-}_{a}-2{\cal D}_{a}F^{+}=$ $0\,$, (5.8d)
$D^{++}P^{(-3)}-{\cal D}^{a}A^{-}_{a}+9S^{--}{\cal J}F^{+}-({\cal J}+2D^{++}S^{--})(M^{-}+N^{-})=$ $0\,$, (5.8e)
$D^{++}\Psi_{\alpha}=$ $0\,$, (5.8f)
$D^{++}\Phi_{\alpha}=$ $0\,$, (5.8g)
$D^{++}\Lambda^{--}_{\alpha}-i{\cal D}_{\alpha\dot{\beta}}\bar{\Phi}^{\dot{\beta}}-({\cal J}+2S^{+-}+2S^{--}D^{++})\Psi_{\alpha}=$ $0\,$, (5.8h)
$D^{++}X^{--}_{\alpha}+i{\cal D}_{\alpha\dot{\beta}}\bar{\Psi}^{\dot{\beta}}-({\cal J}+2S^{+-}+2S^{--}D^{++})\Phi_{\alpha}=$ $0\,.$ (5.8i)
These equations may be solved in terms of the series expansions over the harmonic variables and matching the corresponding powers of harmonics. For instance, Eq. (5.8a) implies $F^{+}(x,u)=f^{i}u^{+}_{i}$; Eqs. (5.8b5.8d) imply $M^{-}(x,u)=M^{i}(x)u^{-}_{i}$, $N^{-}(x,u)=N^{i}(x)u^{-}_{i}$, $A^{-}_{a}(x,u)=A^{i}_{a}(x)u^{-}_{i}$, e.t.c. Note that the fermionic fields $\Psi_{\alpha}$ and $\Phi_{\alpha}$ are harmonic independent, $\Psi_{\alpha}(x,u)=\psi_{\alpha}(x)$, $\Phi_{\alpha}(x,u)=\phi_{\alpha}(x)$. Eliminating all remaining auxiliary fields via their equations of motion and evaluating integrals over harmonic variables, we find the following actions for the physical components:
$S_{q}=\int d^{4}x\,e\left[{\cal D}_{m}\bar{f}_{i}{\cal D}^{m}f^{i}-2S^{2}\bar{f}_{i}f^{i}+\frac{i}{2}\bar{\phi}^{\dot{\beta}}{\cal D}_{\alpha\dot{\beta}}\phi^{\alpha}+\frac{i}{2}\bar{\psi}^{\dot{\beta}}{\cal D}_{\alpha\dot{\beta}}\psi^{\alpha}\right],$ (5.9)
According to Eq. (2.7e), the Riemann curvature tensor in the AdS4 space is $R_{ab}{}^{cd}=-2S^{2}\delta^{c}_{[a}\delta^{d}_{b]}$, and, hence, the Ricci scalar is $R=-12S^{2}$. This allows to match the AdS4 mass term for the scalars in Eq. (5.9) with the one in Ref. [24], $2S^{2}=R/6$.
Acknowledgments. We are grateful to Sergei Kuzenko for proposing to study the problem of analytic action principle in the AdS harmonic superspace and for useful comments. IBS acknowledges the support from the Australian Research Council, project No. DP230101629 during his work at UWA.

### v2

@@ V Example: component structure of the free hypermultiplet action
$D^{++}F^{+}=$ $0\,$, (58a)
$D^{++}M^{-}-({\cal J}+2S^{--}D^{++})F^{+}=$ $0\,$, (58b)
$D^{++}N^{-}-({\cal J}+2S^{--}D^{++})F^{+}=$ $0\,$, (58c)
$D^{++}A^{-}_{a}-2{\cal D}_{a}F^{+}=$ $0\,$, (58d)
$D^{++}\Psi_{\alpha}=0\,,\qquad D^{++}\Phi_{\alpha}=$ $0\,$, (58e)
$D^{++}P^{(-3)}-{\cal D}^{a}A^{-}_{a}+9S^{--}{\cal J}F^{+}-({\cal J}+2D^{++}S^{--})(M^{-}+N^{-})=$ $0\,$, (58f)
$D^{++}\Lambda^{--}_{\alpha}-i{\cal D}_{\alpha\dot{\beta}}\bar{\Phi}^{\dot{\beta}}-({\cal J}+2S^{+-}+2S^{--}D^{++})\Psi_{\alpha}=$ $0\,$, (58g)
$D^{++}X^{--}_{\alpha}+i{\cal D}_{\alpha\dot{\beta}}\bar{\Psi}^{\dot{\beta}}-({\cal J}+2S^{+-}+2S^{--}D^{++})\Phi_{\alpha}=$ $0\,.$ (58h)
Elimination of auxiliary fields from these equations follows the standard procedure [2]: equations (58a)–(58e) are purely kinematic and so may be solved by expanding the series in harmonic variables and comparing term by term, thus eliminating most of the auxiliary fields. Equations (58f)–(58h) represent a combination of kinematic and dynamic equations and allow one to eliminate the rest of the auxiliary fields, and are also responsible for putting the physical fields on shell. For instance, Eq. (58a) implies $F^{+}(x,u)=f^{i}u^{+}_{i}$; Eqs. (58b58d) imply $M^{-}(x,u)=M^{i}(x)u^{-}_{i}$, $N^{-}(x,u)=N^{i}(x)u^{-}_{i}$, $A^{-}_{a}(x,u)=A^{i}_{a}(x)u^{-}_{i}$, e.t.c. Note that the fermionic fields $\Psi_{\alpha}$ and $\Phi_{\alpha}$ are harmonic independent, $\Psi_{\alpha}(x,u)=\psi_{\alpha}(x)$, $\Phi_{\alpha}(x,u)=\phi_{\alpha}(x)$. Eliminating all remaining auxiliary fields via their equations of motion and evaluating integrals over harmonic variables, we find the following actions for the physical components:
$S_{q}=$ $\int d^{4}x\,e\Big[{\cal D}_{m}\bar{f}_{i}{\cal D}^{m}f^{i}-2S^{2}\bar{f}_{i}f^{i}$
$+\frac{i}{2}\bar{\phi}^{\dot{\beta}}{\cal D}_{\alpha\dot{\beta}}\phi^{\alpha}+\frac{i}{2}\bar{\psi}^{\dot{\beta}}{\cal D}_{\alpha\dot{\beta}}\psi^{\alpha}\Big],$ (59)
According to Eq. (7f), the Riemann curvature tensor in the AdS4 space is $R_{ab}{}^{cd}=-2S^{2}\delta^{c}_{[a}\delta^{d}_{b]}$, and, hence, the Ricci scalar is $R=-12S^{2}$. This allows us to match the AdS4 mass term for the scalars in Eq. (59) with the one in Ref. [24], $2S^{2}=R/6$.
In conclusion, it would be interesting to study whether the superfield action (50) can be obtained by dimensional reduction from the superfield theory on AdS5|8 developed in Refs. [22, 23]. In curved space, dimensional reduction is a non-trivial problem and certainly merits future investigation.
Acknowledgments. We are grateful to Sergei Kuzenko for proposing to study the problem of analytic action principle in the AdS4 harmonic superspace and for useful comments. IBS acknowledges the support from the Australian Research Council, project No. DP230101629 during his work at UWA.